# Primordial Black Holes in the Solar System

### v1

@@ Abstract
If primordial black holes (PBHs) of asteroidal mass ($M_{\rm PBH}\in[10^{17},10^{23}]$ g) make up the entire dark matter they could be detectable through their gravitational influence in the solar system. In this work we study the perturbations that PBHs induce on the orbits of planets. Detailed numerical simulations of the solar system, embedded in a halo of primordial black holes, are performed. We show that the perturbations are too small to be directly detectable with current data, challenging recent results that have ruled out PBHs as a dark matter candidate. Using the Earth-Mars distance as an observational probe, we estimate that an improvement in the measurement accuracy by more than an order of magnitude is required to detect the gravitational influence of PBHs in the solar system in the foreseeable future.
@@ 1 Introduction
Recently, Loeb (2024) argued that PBHs within most of the asteroid-mass window are already excluded from making up the entire dark matter, based on solar system ephemerides. The result was obtained by considering the Poissonian fluctuation of the number of PBHs within a given radius $R$ from the Sun. The rate of the change of the total PBH mass, enclosed within $R$, was compared to an observational constraint on the rate of change of the solar mass. This approach assumes that the total mass of PBHs within $R$ has a gravitational effect similar to a point mass and therefore can be added directly to the mass of the Sun. From this, Loeb (2024) concluded that, for $R=50$ AU, PBHs cannot make up the dark matter in the mass range $M_{\rm PBH}\in[6\times 10^{18},10^{22}]$ g. However, Cline (2024) has noted that the choice of $R=50$ AU is not justified and that one cannot easily rule out PBHs as a dark matter candidate.
@@ 2 Methods
We use a second-order Leapfrog integrator with a fixed timestep to simulate the motions of the Sun, the 8 planets and the Earth’s moon. Other solar system objects (SSOs) are neglected as we are only interested in the relative perturbation of planets in the solar system. The goal of this work is not to make accurate predictions on the absolute positions of the bodies of the solar system. Rather, we are interested in the relative perturbation of their position $\delta\mathbb{r}(t)$. We do not expect the smaller bodies to have a sizeable effect on these perturbations. For the same reason we do not treat finite-size effects or relativistic corrections as these will only affect the perturbations at second order. We refer to Tran et al. (2023) for a more detailed discussion on these effects.
The initial conditions for the SSOs are obtained from the Horizons System by JPL, based on the DE441 model (Park et al., 2021). The solar system is embedded in a halo of PBHs filling a box with periodic boundary conditions and length $L_{\rm box}$. The halo is populated with $N_{\rm PBH}=L^{3}_{\rm box}\rho_{\rm CDM}/M_{\rm PBH}$ PBHs of random positions, thus assuming that PBHs make up the entire dark matter with a monochromatic mass distribution. During the simulations, the gravitational forces are only computed for PBHs within a sphere of radius $R_{\rm box}=L_{\rm box}/2$. This makes the comparison with the case of a smooth particle dark matter component more accurate. The PBHs are given a Maxwellian velocity distribution with a dispersion $\sigma_{v}=185$ km/s as well as an additional component due to the rotation of the galactic disk. The latter has a magnitude of $v_{\odot}=230$ km/s, an angle of $60^{\circ}$ w.r.t. the ecliptic plane and a direction such that the relative motion between the PBHs and Earth is maximal on June 1st (Freese et al., 2013). These parameters lead to a relative velocity of the PBHs of $v_{\rm rms}\approx 279$ km/s $\approx$ 59 AU/yr. $\rho_{\rm CDM}=7\times 10^{-25}$ g cm-3 is assumed to facilitate comparisons with Loeb (2024).
@@
$M_{\rm PBH}$ $N_{\rm PBH}$ $R_{\rm box}$ $N_{\rm runs}$ $t$
$10^{18}$ g 18525 200 AU 1000 1 yr
$10^{19}$ g 6252 300 AU 1000 1 yr
$10^{20}$ g 1482 400 AU 3000 1 yr
$10^{21}$ g 500 600 AU 3000 1 yr
$10^{20}$ g 2110 450 AU 500 20 yr
Table 1: Simulation parameters, showing the adopted PBH mass and number, the size of the simulation box, the total number of runs as well as the simulation time.
@@
which has the advantage of being strictly positive, whereas $\delta r/r$ typically oscillates between $-|\delta\mathbb{r}|/r$ and $|\delta\mathbb{r}|/r$ within one synodic orbital period of the SSO considered. While this quantity is more difficult to observe, it will provide an upper bound on $\delta r/r$.
@@
To compare our results with the calculations by Loeb (2024), we perform additional simulations where we approximate the effect of the stochastically changing number of PBHs by corresponding fluctuations of the mass of the Sun. This method was employed in his work where he used a rate
to model the Poisson fluctuations of the number of PBHs within a distance $R$ from the Sun and derive constraints based on that (see Section 4.3 for a detailed discussion of this model). We incorporate this in our simulation setup by adding the PBH halo but instead of computing their gravitational forces we add the mass of all PBHs within $R=50$ AU to the solar mass. This does indeed lead to Poisson fluctuations of the solar mass with the rate given by Equation 4. We compare the results to the standard simulation setup where we treat the gravity that the PBHs exert on the solar system bodies correctly. This enables us to investigate whether the calculations of Loeb (2024) are accurate.
@@ 3 Results
Figure 1 shows the perturbations that are induced on the vector between Earth and Mars for $M_{\rm PBH}\in[10^{18},10^{19},10^{20},10^{21}]$. Due to the random initial conditions of the PBHs for each run there is a significant variance in the value of $|\delta\mathbb{r}|/r$. The median value of the perturbations induced by a halo of PBHs is several orders of magnitude below the detection limit ($\delta r/r\sim 10^{-11}$) in each case and shows no significant dependence on the black hole mass $M_{\rm PBH}$. The mean value of $|\delta\mathbb{r}|/r$ is higher than median and increases slightly with $M_{\rm PBH}$. This can be understood analytically by the impulse approximation which is discussed in Section 4 and that is also displayed in Figure 1. The simulations which treat the effect of the PBHs by a Poisson fluctuation in the solar mass, as was done in the work by Loeb (2024), are shown as a comparison in red. They lead to perturbations which are orders of magnitude higher.
Figure 1: Perturbation of the vector between Earth and Mars, induced by PBHs for different values of their mass $M_{\rm PBH}$. Each blue line corresponds to one simulation run. We only show 100 simulation runs in each case to make comparisons easier and improve readability. The black dashed/dotted line indicates the median/mean value obtained from all simulations. The analytical result of the impulse model for $b_{\rm max}=3b_{\rm min}$ is shown as an orange line. The results from 10 simulation runs which treat the effects of the PBHs by Poisson fluctuations of the solar mass are shown in red. We have excluded them for $M_{\rm PBH}=10^{21}$ g as they exceed $|\delta\mathbb{r}|/r=10^{-12}$ in that case.
@@
Figure 2: Left: Probability distribution function of the perturbation strength $|\delta\mathbb{r}|/r$, evaluated at $t=1.0$ yr for different values of the PBH mass $M_{\rm PBH}$. The black dashed line indicates a power law slope of -2. Vertical, dashed lines indicate the critical perturbation $|\delta\mathbb{r}|/r_{\rm crit}$, that is defined in Equation 12, for each value of $M_{\rm PBH}$. Right: Perturbation strength $|\delta\mathbb{r}|/r$, evaluated at $t=1.0$ yr vs. the minimum impact parameter $b_{\rm min}$ observed within the simulation period. The black, dashed lines correspond to Equation 8. The dotted line for $M_{\rm PBH}=10^{21}$ g takes into account the finite distance travelled by the PBHs within the simulation time and is explained in Section 4.5.
@@
In the right panel of Figure 3 we show the mean perturbation strength $|\delta\mathbb{r}|/r$ of all simulation runs for several planets. In our analysis we focus on the Earth-Mars pair for the reasons outlined in Section 2. The results shown here demonstrate that this does not limit the generality of our statements as the observed perturbations are very similar for other planets in the solar system.
Figure 3: Left: Perturbation of the vector between Earth and Mars induced by PBHs for $M_{\rm PBH}=10^{20}$ g as a function of time. Each blue line corresponds to one simulation run. We only show 100 simulation runs to improve readability. The black dashed/dotted line indicates the median/mean value obtained from all 500 simulations. The analytical result of the impulse model is displayed as an orange line. The green line shows a model for the median perturbation strength, given by Equation 15. The red line is an estimate of the $3\sigma$ observational detection limit. Right: Mean perturbation strength $|\delta\mathbb{r}|/r$ as a function of time for the vector between Earth and various planets.
@@ 4.1 Impulse Model
So far, we have considered the perturbation from a single encounter with a PBH. For a DM halo composed of PBHs the rate of encounters in the solar system will be $\Gamma=\sigma/m\rho v=\pi b^{2}/M_{\rm PBH}\rho_{\rm CDM}v_{\rm PBH}$. This means that the number of scattering events per interval of $b$ is $\mathrm{d}N/\mathrm{d}b=2\pi b/M_{\rm PBH}\rho_{\rm CDM}v_{\rm PBH}t$. From this we can obtain the perturbation strength,
@@
The fact that the perturbations are dominated by the closest encounter is also demonstrated in the right panel of Figure 2, where we plot the perturbation strength $|\delta\mathbb{r}|/r$ at $t=1$ yr against the smallest impact parameter within the simulation run. The observed values agree reasonably well with Equation 8 that is displayed as dashed lines. Note that here we have assumed a singular encounter with a PBH that moves perpendicular to the ecliptic plane. Therefore some deviation from the analytical estimates is expected in the more complex simulation environment. In addition, the approximation of infinite integration boundaries in Equation 5 breaks down for large PBH masses which explains the offset observed for $M_{\rm PBH}=10^{21}$ g. This will be discussed in more detail in Section 4.4.
Equation 9 provides a simple explanation for the observed weak dependence of the perturbation strength $|\delta\mathbb{r}|/r$ on the black hole mass as it predicts only a logarithmic dependence as long as $b_{\rm max}\gg r$. The underlying reason is that the strength of the gravitational force ($\sim M_{\rm PBH}$) is counteracted by the decrease in the scattering rate ($\sim 1/M_{\rm PBH}$). This means, our simulation results should have some validity beyond the mass range considered here. However, if PBHs are sufficiently light (or $t$ sufficiently long) then one could have $b_{\rm max}\ll r$ (assuming again that $b_{\rm max}\sim\mathcal{O}(b_{\rm min})$ as we argued above) and thus $|\delta\mathbb{r}|/r\sim\log(1+b_{\rm max}/r)\approx b_{\rm max}/r\sim\sqrt{M_{\rm PBH}}$. For $t=1$ yr one has $b_{\rm min}=r$ for $M_{\rm PBH}\sim 10^{18}$ g. Indeed, in our simulation results we notice a decreasing strength of the mean amplitude of the perturbations as $M_{\rm PBH}$ decreases (compare the dotted black lines across the four panels in Figure 1). This also provides consistency as the perturbations should disappear as $M_{\rm PBH}\rightarrow 0$ and one obtains a smooth DM component.
@@ 4.2 Statistics of the perturbations
The values of $|\delta\mathbb{r}|/r_{\rm crit}$ are displayed as vertical lines in the left panel of Figure 2. For $|\delta\mathbb{r}|/r>|\delta\mathbb{r}|/r_{\rm crit}$ the approximation $b\gg r$ breaks down. For $M_{\rm PBH}\in[10^{20},10^{21}]$ g most of the observed perturbations are below the critical value and indeed $p(|\delta\mathbb{r}|/r)$ follows a power law with a slope of -2 to good agreement. For $M_{\rm PBH}=10^{18}$ g, most of the perturbations are far above the critical value and our approximation breaks down, which explains the observed deviation from the expected power law. At $M_{\rm PBH}=10^{19}$ we observe a power law with a slope of -2 for perturbations up to $|\delta\mathbb{r}|/r\sim 10|\delta\mathbb{r}|/r_{\rm crit}$ which suggests that Equation 11 is valid up to that point. We cannot make strong statements for the behaviour of the distribution function beyond this, based on our simulations data, as the occurrence of such perturbations is too rare and we are thus limited by the number of simulations that we have performed. If we trust the analytical model from Section 4.1 for $b\ll r$ then, using Equation 7, one would expect $p(|\delta\mathbb{r}|/r)\sim(|\delta\mathbb{r})^{-3}$ in this regime.
4.3 Poissonian Fluctuations and the Failure of the Central Point Mass Approximation
We briefly want to discuss the role of Poissonian fluctuations with regards to the effects of PBHs on solar system ephemerides. If the dark matter is composed of asteroid-mass black holes then their number within the solar system will undergo sizeable Poisson fluctuations. Loeb (2024) considered the rate of change in mass enclosed within a radius $R$, as given by Equation 4. In his work it is then assumed that such fluctuations would induce similar effects as if a point mass in the center of the solar system would change its mass with the same rate, effectively changing the solar mass by a small amount. For $R=50$ AU and $M_{\rm PBH}\in[6\times 10^{18},10^{22}]$ g, this would lead to a rate larger than the constraint on $\dot{M}_{\rm sol}$ by Pitjeva et al. (2021) and thus PBHs in this mass range cannot make up the entire dark matter. However, the assumption that the Poisson fluctuations of the mass of PBHs enclosed within a given radius $R$ can be equated with a change in the solar mass is not justified. The main reason is that the fluctuations of the PBH density within the solar system are random and do not follow spherical symmetry. Figure 1 demonstrates that the simulations which treat the gravitational influence of PBHs by Poission fluctuations of the solar mass predict perturbations that are orders of magnitude larger than actual perturbations, found when the PBHs are treated as individual objects. One therefore can not equate the random fluctuations of the PBH density in the solar system with a change in the mass of the Sun.

### v2

@@ Abstract
If primordial black holes (PBHs) of asteroidal mass make up the entire dark matter they could be detectable through their gravitational influence in the solar system. In this work, we study the perturbations that PBHs induce on the orbits of planets. Detailed numerical simulations of the solar system, embedded in a halo of PBHs, are performed. Using the Earth-Mars distance as an observational probe, we show that the perturbations are below the current detection limits and thus PBHs are not directly constrained by solar system ephemerides. We estimate that an improvement in the measurement accuracy by more than an order of magnitude or the extraction of signals well below the noise level are required to detect the gravitational influence of PBHs in the solar system in the foreseeable future.
software: Julia (Bezanson et al., 2017), Matplotlib (Hunter, 2007)
@@ 1 Introduction
Recently, Loeb (2024) argued that PBHs within most of the asteroid-mass window are already excluded from making up the entire dark matter, based on their perturbations of solar system bodies. The result was obtained by considering the Poissonian fluctuation of the number of PBHs within a given radius $R$ from the Sun. The rate of the change of the total PBH mass, enclosed within $R$, was compared to an observational constraint on the rate of change of the solar mass. This approach assumes that the total mass of PBHs within $R$ has a gravitational effect similar to a point mass and therefore can be added directly to the mass of the Sun. From this, Loeb (2024) concluded that, for $R=50$ AU, PBHs cannot make up the entire dark matter in the mass range $M_{\rm PBH}\in[6\times 10^{18},10^{22}]$ g. However, Cline (2024) has noted that the choice of $R=50$ AU is not justified and that one cannot easily rule out PBHs as a dark matter candidate.
@@ 2 Methods
We use a second-order Leapfrog integrator with a fixed timestep to simulate the motions of the Sun, the 8 planets, and the Earth’s moon. Other solar system objects (SSOs) are neglected as we are only interested in the relative perturbation of planets in the solar system. The goal of this work is not to make accurate predictions on the absolute positions of the bodies of the solar system. Rather, we are interested in the relative perturbation of their position $\delta\mathbb{r}(t)$. We do not expect the smaller bodies to have a sizeable effect on these perturbations. For the same reason we do not treat finite-size effects or relativistic corrections as these will only affect the perturbations at second order. We refer to Tran et al. (2023) for a more detailed discussion on these effects.
The initial conditions for the SSOs are obtained from the Horizons System by JPL, based on the DE441 model (Park et al., 2021). The solar system is embedded in a halo of PBHs filling a box with periodic boundary conditions and length $L_{\rm box}$. The halo is populated with $N_{\rm PBH}=L^{3}_{\rm box}\rho_{\rm CDM}/M_{\rm PBH}$ PBHs of random positions, thus assuming that PBHs make up the entire dark matter with a monochromatic mass distribution. During the simulations, the gravitational forces are only computed for PBHs within a sphere of radius $R_{\rm box}=L_{\rm box}/2$. This makes the comparison with the case of a smooth particle dark matter component more accurate. The PBHs are given a Maxwellian velocity distribution with a dispersion $\sigma_{v}=185$ km s-1 as well as an additional component due to the rotation of the galactic disk. The latter has a magnitude of $v_{\odot}=230$ km s-1, an angle of $60^{\circ}$ w.r.t. the ecliptic plane and a direction such that the relative motion between the PBHs and Earth is maximal on June 1st (Freese et al., 2013). These parameters lead to a relative velocity of the PBHs of $v_{\rm rms}\approx 279$ km s-1 $\approx 59$ AU yr-1. $\rho_{\rm CDM}=7\times 10^{-25}$ g cm-3 is assumed to facilitate comparisons with Loeb (2024).
@@
$M_{\rm PBH}$ (g) $N_{\rm PBH}$ $R_{\rm box}$ (AU) $N_{\rm runs}$ $t$ (yr)
$10^{18}$ 18525 200 1000 1
$10^{19}$ 6252 300 1000 1
$10^{20}$ 1482 400 3000 1
$10^{21}$ 500 600 3000 1
$10^{20}$ 2110 450 500 20
Table 1: Simulation parameters, showing the adopted PBH mass and number, the size of the simulation box, the total number of runs, as well as the simulation time.
@@
which has the advantage of being strictly positive, whereas $\delta r/r$ typically oscillates between $-|\delta\mathbb{r}|/r$ and $|\delta\mathbb{r}|/r$ within one synodic orbital period of the SSO considered. While this quantity is more difficult to observe, it provides a reasonable estimate of the largest observable perturbation $\delta r/r$ per orbital period, as we show in Section 3.
@@ 3 Results
Figure 1 shows the perturbations that are induced on the vector between Earth and Mars for $M_{\rm PBH}\in[10^{18},10^{19},10^{20},10^{21}]$ g. Due to the random initial conditions of the PBHs for each run there is a significant variance in the value of $|\delta\mathbb{r}|/r$. The median value of the perturbations induced by a halo of PBHs is several orders of magnitude below the detection limit ($\delta r/r\sim 10^{-11}$) in each case and shows no significant dependence on the black hole mass $M_{\rm PBH}$. The mean value of $|\delta\mathbb{r}|/r$ is higher than median and increases slightly with $M_{\rm PBH}$. This can be understood analytically by the impulse approximation which is discussed in Section 4 and that is also displayed in Figure 1.
Figure 1: Perturbation of the vector between Earth and Mars, induced by PBHs for different values of their mass $M_{\rm PBH}$. Each blue line corresponds to one simulation run. We only show 100 simulation runs in each case to make comparisons easier and improve readability. The black dashed/dotted line indicates the median/mean value obtained from all simulations. The analytical result of the impulse model for $b_{\rm max}=3b_{\rm min}$ is shown as an orange line.
@@
Figure 2: Left: Probability distribution function of the perturbation $|\delta\mathbb{r}|/r$ of the Earth-Mars vector, evaluated at $t=1.0$ yr for different values of the PBH mass $M_{\rm PBH}$. The black dashed line indicates a power law with slope of -2. Vertical, dashed lines indicate the critical perturbation $\left(|\delta\mathbb{r}|/r\right)_{\rm crit}$, that is defined in Equation 11, for each value of $M_{\rm PBH}$. Right: Perturbation strength $|\delta\mathbb{r}|/r$ of the Earth-Mars vector, evaluated at $t=1.0$ yr vs. the minimum impact parameter $b_{\rm min}$ observed within the simulation period. The black, dashed lines correspond to Equation 7. The dotted line for $M_{\rm PBH}=10^{21}$ g takes into account the finite distance travelled by the PBHs within the simulation time and is explained in Section 4.4.
@@
In the right panel of Figure 3 we show the mean perturbation strength $|\delta\mathbb{r}|/r$ of all simulation runs for the vector between Earth and each planet. Notably, $|\delta\mathbb{r}|/r$ is very similar for the inner planets while it decreases for the outer planets as the distance to the Earth gets larger. This can be understood by the impulse model (see Section 4.1) which predicts two different regimes, depending on whether the typical impact parameter of the PBH $b$ is much smaller or much larger than the distance $r$ between the Earth and the other SSO. Since $b_{\rm min}\approx 3.4$ AU for our parameters (from Equation 9) one has $b_{\rm min}>r$ for the inner planets and $b_{\rm min}<r$ for the outer planets.
In the results presented in this study, we analyse the perturbation of the vector $|\delta\mathbb{r}|$ between the Earth and other SSOs. However, only the perturbation of the distance $\delta r$ can be observed with high accuracy. We argue that this should not affect our results as $\delta r$ typically oscillates between $-|\delta\mathbb{r}|$ and $|\delta\mathbb{r}|$ and thus our results can be regarded as an upper bound of the observable perturbation since $|\delta r|\leq|\delta\mathbb{r}|$ by definition. To test this statement we compare the values of $\delta r$ and $\delta\mathbb{r}$ for the vector between Earth and Mars. We use the data from the last synodic orbital period from the 500 simulations performed over a timespan of $t=20$ yr. We find that 87% of the simulations reach $|\delta r|>|\delta\mathbb{r}|/2$ at least once and 72% of the simulations reach $|\delta r|>|\delta\mathbb{r}|/2$ during at least 75% of the respective timespan. In addition, 50% of the simulations reach $|\delta r|>0.9|\delta\mathbb{r}|$ at least once within the orbital period. This justifies our approach of studying the perturbation of the vector $\delta\mathbb{r}$ as a good approximation of the largest observable perturbation of the distance $\delta r$.
Figure 3: Left: Perturbation of the vector between Earth and Mars induced by PBHs for $M_{\rm PBH}=10^{20}$ g as a function of time. Each blue line corresponds to one simulation run. We only show 100 simulation runs to improve readability. The black dashed/dotted line indicates the median/mean value obtained from all 500 simulations. The analytical result of the impulse model is displayed as an orange line. The green line shows a model for the median perturbation strength, given by Equation 14. The red line is an estimate of the $3\sigma$ observational detection limit. Right: Mean perturbation strength $|\delta\mathbb{r}|/r$ as a function of time for the vector between Earth and each planet.
@@ 4.1 Impulse Model
So far, we have considered the perturbation from a single encounter with a PBH. For a DM halo composed of PBHs the rate of encounters in the solar system will be $\Gamma=(\sigma/m)\rho v=(\pi b^{2}/M_{\rm PBH})\rho_{\rm CDM}v_{\rm PBH}$. This means that the number of scattering events per interval of $b$ is $\mathrm{d}N/\mathrm{d}b=(2\pi b/M_{\rm PBH})\rho_{\rm CDM}v_{\rm PBH}t$. From this we can obtain the perturbation strength,
@@
The fact that the perturbations are dominated by the closest encounter is also demonstrated in the right panel of Figure 2, where we plot the perturbation strength $|\delta\mathbb{r}|/r$ at $t=1$ yr against the smallest impact parameter within the simulation run. The observed values agree reasonably well with Equation 7 that is displayed as dashed lines. Note that here we have assumed a singular encounter with a PBH that moves perpendicular to the ecliptic plane. Therefore, some deviation from the analytical estimates is expected in the more complex simulation environment. In addition, the approximation of infinite integration boundaries in Equation 4 breaks down for large PBH masses which explains part of the offset observed for $M_{\rm PBH}=10^{21}$ g. This will be discussed in more detail in Section 4.3.
Equation 8 provides a simple explanation for the observed weak dependence of the perturbation strength $|\delta\mathbb{r}|/r$ on the black hole mass as it predicts only a logarithmic dependence as long as $b_{\rm max}\gg r$. The underlying reason is that the strength of the gravitational force ($\sim M_{\rm PBH}$) is counteracted by the decrease in the scattering rate ($\sim 1/M_{\rm PBH}$). This means, our simulation results should have some validity beyond the mass range considered here. However, if PBHs are sufficiently light (or $t$ sufficiently long) then one could have $b_{\rm max}\ll r$ (assuming again that $b_{\rm max}\sim\mathcal{O}(b_{\rm min})$ as we argued above) and thus $|\delta\mathbb{r}|/r\sim\log(1+b_{\rm max}/r)\approx b_{\rm max}/r\sim\sqrt{M_{\rm PBH}}$. For $t=1$ yr one has $b_{\rm min}=r$ for $M_{\rm PBH}\sim 10^{18}$ g. Indeed, in our simulation results we notice a decreasing strength of the mean amplitude of the perturbations as $M_{\rm PBH}$ decreases (compare the dotted black lines across the four panels in Figure 1). In addition, this explains the declining strength of $|\delta\mathbb{r}|/r$ for the outer planets in the right panel of Figure 3 as one has $b_{\rm max}<r$ there. Finally, we want to emphasize that the decreasing strength of the perturbations at low PBH masses provides consistency as they should disappear as $M_{\rm PBH}\rightarrow 0$ and one obtains a smooth DM component.
@@ 4.2 Statistics of the perturbations
The values of $|\delta\mathbb{r}|/r_{\rm crit}$ are displayed as vertical lines in the left panel of Figure 2. For $|\delta\mathbb{r}|/r>|\delta\mathbb{r}|/r_{\rm crit}$ the approximation $b\gg r$ breaks down. For $M_{\rm PBH}\in[10^{20},10^{21}]$ g most of the observed perturbations are below the critical value and indeed $p(|\delta\mathbb{r}|/r)$ follows a power law with a slope of -2 to good agreement. For $M_{\rm PBH}=10^{18}$ g, most of the perturbations are far above the critical value and our approximation breaks down, which explains the observed deviation from the expected power law. At $M_{\rm PBH}=10^{19}$ we observe a power law with a slope of -2 for perturbations up to $|\delta\mathbb{r}|/r\sim 10|\delta\mathbb{r}|/r_{\rm crit}$ which suggests that Equation 10 is valid up to that point. We cannot make strong statements for the behaviour of the distribution function beyond this, based on our simulation data, as the occurrence of such perturbations is too rare and we are thus limited by the number of simulations that we have performed. If we trust the analytical model from Section 4.1 for $b\ll r$ then, using Equation 6, one would expect $p(|\delta\mathbb{r}|/r)\sim(|\delta\mathbb{r})^{-3}$ in this regime.